OpenStudy (anonymous):

Pleassse help! medal will be given Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (imstuck):

The first thing you need to do is to find the slope of the line using the given points R and S. Then you have to put it into Ax + By = C format, which is standard form for a line, with the A being the slope, and the C being the y intercept. Do you know how to find the slope?

OpenStudy (anonymous):

i think i do, isnt it y1-y2/ x1-x2?

OpenStudy (imstuck):

yes

OpenStudy (imstuck):

What do you get the slope to be, then?

OpenStudy (imstuck):

Actually, its y2 - y1/x2 - x1.

OpenStudy (anonymous):

-18?

OpenStudy (anonymous):

@IMStuck

OpenStudy (princeharryyy):

(y-3)=((-6-3)/(-6-3))*(x-3); y-3=x-3; =>y=x

OpenStudy (princeharryyy):

@Jessica4321 ????